# Exponentially Enhanced Scheme for the Heralded Qudit GHZ State in Linear Optics

### v1

@@ Abstract
High-dimensional multipartite entanglement plays a crucial role in quantum information science. However, existing schemes for generating such entanglement become increasingly complex and costly as the dimension of quantum units increases. In this work, we overcome the limitation by proposing a significantly enhanced linear optical heralded scheme that generates the $d$-level $N$-partite GHZ state with single-photon sources and their linear operations. Our scheme requires $dN$ photons to generate the target state with substantially improved success probability from previous schemes. It employs linear optical logic gates compatible with any qudit encoding system and can generate generalized GHZ states with installments of beamsplitters. With efficient generations of high-dimensional resource states, our work opens avenues for further exploration in high-dimensional quantum information processing.
@@
Numerous advantages arise from entangling multiple quantum units in quantum information processing. The Bell state of qubits facilitates quantum enhancements over classical counterparts [4, 5]. Beyond two-qubit systems, quantum entanglement in multipartite and/or higher-dimensional systems is essential for new quantum technologies such as quantum cryptography, reduction of communication complexity, quantum communication protocols, and random number generation [6, 7, 8, 9, 10].
@@
We contribute to overcoming this limit by providing a magnificently enhanced heralded scheme for generating the qudit multipartite entanglement in linear optical networks (LONs). The LON is one of promising systems that can feasibly generate quantum resource states. Non-destructive entanglement generations by heralding are possible in LONs by utilizing ancillary particles to herald the successful operation events without direct detections of the target states. Therefore, photons entangled by heralding become proper resources for quantum computations with multiple gates [14, 15] and also facilitates loophole-free Bell tests [16, 17]. However, the number of additional particles and modes that the heralded schemes require increases as the dimension of entanglement becomes higher, which makes the circuit structures much more intricate. Therefore, it is prerequisite to design optimized circuits that achieve the target state with minimal resources for the efficient high-dimensional quantum information processing.
@@
where $|j\rangle_{k}$ denotes the $k$th qudit being in the $j$th state. The qudit GHZ state reveals intriguing non-classical properties of $N$-partite $d$-level systems such as the GHZ paradox [18, 19, 20, 21] and Bell inequalities [22]. It also becomes a resource state to generate qudit cluster states for the high-dimensional measurement-based quantum computing (MBQC) [23, 24, 25].
Heralded schemes for the qudit GHZ state in various physical setups have been proposed, e.g., with single-photon sources and linear optical operators [25], EPR sources [26], or an array of non-interacting photon emitters [27]. Among them, Ref. [25] suggested a linear optical heralded scheme based on the zero-transmission laws (ZTL) in quantum Fourier tranforms (QFTs). However, it suffers from low efficiency and exponentially increasing photon numbers as $N$ and $d$ grows.
In this work, we propose a scheme that generates the GHZ state of any $N$ and $d$ with the most feasible optical elements, i.e., single photons and linear optical operators as in Ref. [25]. Our scheme has several manifest advantages over formerly suggested schemes: First, it is an efficient circuit with significantly less photons ($dN$ photons for $|GHZ_{N}^{d}\rangle$) and exponentially higher success probability than that in Ref. [25]. Second, it can easily control the probability amplitudes of each term in the GHZ state by simply placing beamsplitters on the paths of photons, which result in the generation of the generalized qudit GHZ state. Third, it consists of linear optical logic gates that any type of qubit encoding (e.g., optical angular momentum (OAM), multi-rail, time-bin, etc.) can establish.
@@
1. $CN_{d}$: The $d$-level generalized CNOT gate between the spatial mode $j$ ($(\in\{0,1,\cdots,d\})$ and the internal state $\tilde{s}$ ($(\in\{\tilde{0},\tilde{1},\cdots,\tilde{d}\})$, which works on states as
$|\tilde{s}\rangle_{j}={\hat{a}}^{\dagger}_{j,\tilde{s}}|vac\rangle$
$\to$ $~{}CN_{d}|\tilde{s}\rangle_{j}=|\tilde{s}\rangle_{j\oplus_{d}\tilde{s}}={\hat{a}}^{\dagger}_{j\oplus_{d}\tilde{s},\tilde{s}}|vac\rangle$ (2)
(${\hat{a}^{\dagger}}_{j,\tilde{s}}$ is a photon creation operator, $|vac\rangle$ is the vacuum state, and $\oplus_{d}$ is the addition modular $d$). Its inverse $\overline{CN_{d}}$ works as
$|\tilde{s}\rangle_{j}={\hat{a}}^{\dagger}_{j,\tilde{s}}|vac\rangle$
$\to$ $~{}\overline{CN}_{d}|\tilde{s}\rangle_{j}=|\tilde{s}\rangle_{j\ominus_{d}\tilde{s}}={\hat{a}}^{\dagger}_{j\ominus_{d}\tilde{s},\tilde{s}}|vac\rangle$ (3)
@@ 2. $F_{d}$: The internal state Fourier-transformation gate,
$F_{d}|\tilde{s}\rangle_{j}=|s\rangle_{j}=\frac{1}{\sqrt{d}}\sum_{r=0}^{d-1}e^{i\omega r}|\tilde{r}\rangle~{}~{}~{}(\omega=e^{i\frac{2\pi k}{d}}).$ (4)
@@ 3. $X_{d}$: The $d$-level unitary generalization of $\sigma_{x}$ in qubit,
4. $BS_{t}$: A beamsplitter that transmits a photon with transitivity $t$.
5. $D_{F}$: Number resolving detections in the Fourier-transformed basis, i.e., measurements in $\{|s\rangle\}_{s=0}^{d-1}$ and postselections of the cases when only one particle arrives at the detector. This is achieved by combining $F_{d}$ and $CN_{d}$, i.e.,
@@
By combining the gates listed in Fig. 1, we can design a circuit that generates $|GHZ_{N}^{d}\rangle$ for any $N$ and $d$. Our scheme is given in Fig. 2, which can be considered a linear optical installation of the boson subtraction scheme in Ref. [28] to generate the same target state. While Ref. [28] lacks how to furnish the qudit boson subtraction operators with real physical systems, we demonstrate that the elements of the boson subtraction scheme for $|GHZ_{N}^{d}\rangle$ can be translated to the linear optical logic gates to construct a physical qudit GHZ scheme with heralding. A more detailed explanation on the relation between the subtraction operator and our optical heralded scheme is given in Appendix B.
In each input mode of the circuit in Fig. 2, we inject $d$ photons with different internal states in the same mode. The operation of each dashed box in the circuit corresponds to the boson subtraction of $(d-1)$ particles, hence $(d-1)N$ subtraction of photons in total. $BS_{t}$s are attached to control the probability amplitude of the target state, hence we can obtain different states in the class of generalized GHZ states by changing the transitivity of the $BS_{t}$s. It is worth noting that our circuit inherits the cyclic symmetry of the qudit GHZ state.
@@
Figure 2: The linear optical circuit that generates $|GHZ_{N}^{d}\rangle.$ ABSs are attached at the end of the $d$ sub-modes of the first spatial subsystem $1$ to control the probability amplitudes of the terms in the target state, hence we can obtain the generalized qudit GHZ states.
@@
As proof of concept, we provide here a step-by-step analysis on the $N=d=3$ example (Fig. 3). There are three spatial subsystems $\{a,b,c\}$ that has three sub-modes respectively, hence the photon creation operator is denoted as $\{{\hat{a}^{\dagger}}_{j,s},{\hat{b}^{\dagger}}_{k,r},{\hat{c}^{\dagger}}_{l,t}\}$ where $\{j,k,l\}$ denotes the spatial sub-modes and $\{s,r,t\}$ the qudit states ($j,k,l,s,r,t\in\{0,1,2\}$).
@@
Figure 3: $|GHZ_{3}^{3}\rangle$ example of our scheme. We need a BS with transitivity $\frac{1}{\sqrt{3}}$ on $X$, $Y$ and $Z$ to carve the extra amplitude of $|111\rangle$.
@@
Step 1. Initial state preparation: We inject three photons of orthogonal internal states with each other into the 0th modes of $a$, $b$, and $c$ respectively, hence the initial state is prepared as
${\hat{a}}^{\dagger}_{0,0}{\hat{a}}^{\dagger}_{0,1}{\hat{a}}^{\dagger}_{0,2}{\hat{b}}^{\dagger}_{0,0}{\hat{b}}^{\dagger}_{0,1}{\hat{b}}^{\dagger}_{0,2}\hat{c}^{\dagger}_{0,0}\hat{c}^{\dagger}_{0,1}\hat{c}^{\dagger}_{0,2}|vac\rangle.$ (7)
@@
Step 2. $CN_{3}$ operation: This operator works on the basis of $\{|\tilde{j}\rangle\}_{j=0}^{2}$, hence the initial state is transformed as
$\frac{1}{\sqrt{3}^{9}}$ $\big{(}{\hat{a}}^{\dagger 3}_{0,\tilde{0}}+{\hat{a}}^{\dagger 3}_{1,\tilde{1}}+{\hat{a}}^{\dagger 3}_{2,\tilde{2}}-3{\hat{a}}^{\dagger}_{0,\tilde{0}}{\hat{a}}^{\dagger}_{1,\tilde{1}}{\hat{a}}^{\dagger}_{2,\tilde{2}}\big{)}$
×$\times\big{(}{\hat{b}}^{\dagger 3}_{0,\tilde{0}}+{\hat{b}}^{\dagger 3}_{1,\tilde{1}}+{\hat{b}}^{\dagger 3}_{2,\tilde{2}}-3{\hat{b}}^{\dagger}_{0,\tilde{0}}{\hat{b}}^{\dagger}_{1,\tilde{1}}{\hat{b}}^{\dagger}_{2,\tilde{2}}\big{)}$
×$\times\big{(}\hat{c}^{\dagger 3}_{0,\tilde{0}}+\hat{c}^{\dagger 3}_{1,\tilde{1}}+\hat{c}^{\dagger 3}_{2,\tilde{2}}-3\hat{c}^{\dagger}_{0,\tilde{0}}\hat{c}^{\dagger}_{1,\tilde{1}}\hat{c}^{\dagger}_{2,\tilde{2}}\big{)}|vac\rangle$ (8)
@@
Step 3. Rewiring: Now we rewire so that creation operators $\{{\hat{a}^{\dagger}}_{1,\tilde{1}},{\hat{b}^{\dagger}}_{1,\tilde{1}},{\hat{c}^{\dagger}}_{1,\tilde{1}}\}$ are sent to $\{\hat{X}^{\dagger}_{\tilde{1}},\hat{Y}^{\dagger}_{\tilde{1}},\hat{Z}^{\dagger}_{\tilde{1}}\}$ and $\{{\hat{a}^{\dagger}}_{2,\tilde{2}},{\hat{b}^{\dagger}}_{2,\tilde{2}},{\hat{c}^{\dagger}}_{2,\tilde{2}}\}$ to $\{{\hat{b}^{\dagger}}_{2,\tilde{2}},{\hat{c}^{\dagger}}_{2,\tilde{2}},{\hat{a}^{\dagger}}_{2,\tilde{2}}\}$ :
$\frac{1}{\sqrt{3}^{9}}$ $\big{(}{\hat{a}}^{\dagger 3}_{0,\tilde{0}}+\hat{X}^{\dagger 3}_{\tilde{1}}+{\hat{b}}^{\dagger 3}_{2,\tilde{2}}-3{\hat{a}}^{\dagger}_{0,\tilde{0}}\hat{X}^{\dagger}_{\tilde{1}}{\hat{b}}^{\dagger}_{2,\tilde{2}}\big{)}$
×$\times\big{(}{\hat{b}}^{\dagger 3}_{0,\tilde{0}}+\hat{Y}^{\dagger 3}_{\tilde{1}}+\hat{c}^{\dagger 3}_{2,\tilde{2}}-3{\hat{b}}^{\dagger}_{0,\tilde{0}}\hat{Z}^{\dagger}_{\tilde{1}}\hat{c}^{\dagger}_{2,\tilde{2}}\big{)}$
×$\times\big{(}\hat{c}^{\dagger 3}_{0,\tilde{0}}+\hat{Z}^{\dagger 3}_{\tilde{1}}+\hat{a}^{\dagger 3}_{2,\tilde{2}}-3\hat{c}^{\dagger}_{0,\tilde{0}}\hat{X}^{\dagger}_{\tilde{1}}\hat{a}^{\dagger}_{2,\tilde{2}}\big{)}|vac\rangle$ (9)
Step 4. $X_{d}$ and $F_{3}$: Now the internal states of the creation operators go back to the basis of $\{|j\rangle\}_{j=0}^{2}$.
$\frac{1}{\sqrt{3}^{9}}$ $\big{(}{\hat{a}}^{\dagger 3}_{0,0}+\hat{X}^{\dagger 3}_{\tilde{1}}+{\hat{b}}^{\dagger 3}_{2,0}-3{\hat{a}}^{\dagger}_{0,0}\hat{X}^{\dagger}_{\tilde{1}}{\hat{b}}^{\dagger}_{2,0}\big{)}$
$\times$ $\big{(}{\hat{b}}^{\dagger 3}_{0,0}+\hat{Y}^{\dagger 3}_{\tilde{1}}+{\hat{c}}^{\dagger 3}_{2,0}-3{\hat{b}}^{\dagger}_{0,0}\hat{Y}^{\dagger}_{\tilde{1}}{\hat{c}}^{\dagger}_{2,0}\big{)}$
$\times$ $\big{(}\hat{c}^{\dagger 3}_{0,0}+\hat{Z}^{\dagger 3}_{\tilde{1}}+{\hat{a}}^{\dagger 3}_{2,0}-3\hat{c}^{\dagger}_{0,0}\hat{Z}^{\dagger}_{\tilde{1}}{\hat{a}}^{\dagger}_{2,0}\big{)}|vac\rangle$ (10)
@@
$\frac{1}{3^{9}}\Big{[}({\hat{a}^{\dagger}}_{0,\tilde{0}}+{\hat{a}^{\dagger}}_{2,\tilde{1}}+{\hat{a}^{\dagger}}_{1,\tilde{2}})^{3}+\hat{X}^{\dagger 3}_{\tilde{1}}+({\hat{b}^{\dagger}}_{2,\tilde{0}}+{\hat{b}^{\dagger}}_{1,\tilde{1}}+{\hat{b}^{\dagger}}_{0,\tilde{2}})^{3}$
$~{}~{}~{}-3\sqrt{3}\alpha({\hat{a}^{\dagger}}_{0,\tilde{0}}+{\hat{a}^{\dagger}}_{2,\tilde{1}}+{\hat{a}^{\dagger}}_{1,\tilde{2}})\hat{X}^{\dagger}_{\tilde{1}}({\hat{b}^{\dagger}}_{2,\tilde{0}}+{\hat{b}^{\dagger}}_{1,\tilde{1}}+{\hat{b}^{\dagger}}_{0,\tilde{2}})\Big{]}$
×$~{}\times\Big{[}({\hat{b}^{\dagger}}_{0,\tilde{0}}+{\hat{b}^{\dagger}}_{2,\tilde{1}}+{\hat{b}^{\dagger}}_{1,\tilde{2}})^{3}+\hat{Y}^{\dagger 3}_{\tilde{1}}+({\hat{c}^{\dagger}}_{2,\tilde{0}}+{\hat{c}^{\dagger}}_{1,\tilde{1}}+{\hat{c}^{\dagger}}_{0,\tilde{2}})^{3}$
$~{}~{}~{}-3\sqrt{3}\alpha({\hat{b}^{\dagger}}_{0,\tilde{0}}+{\hat{b}^{\dagger}}_{2,\tilde{1}}+{\hat{b}^{\dagger}}_{1,\tilde{2}})\hat{Y}^{\dagger}_{\tilde{1}}({\hat{c}^{\dagger}}_{2,\tilde{0}}+{\hat{c}^{\dagger}}_{1,\tilde{1}}+{\hat{c}^{\dagger}}_{0,\tilde{2}})\Big{]}$
×$~{}\times\Big{[}({\hat{c}^{\dagger}}_{0,\tilde{0}}+{\hat{c}^{\dagger}}_{2,\tilde{1}}+{\hat{c}^{\dagger}}_{1,\tilde{2}})^{3}+\hat{Z}^{\dagger 3}_{\tilde{1}}+({\hat{a}^{\dagger}}_{2,\tilde{0}}+{\hat{a}^{\dagger}}_{1,\tilde{1}}+{\hat{a}^{\dagger}}_{0,\tilde{2}})^{3}$
$~{}~{}~{}-3\sqrt{3}\alpha({\hat{c}^{\dagger}}_{0,\tilde{0}}+{\hat{c}^{\dagger}}_{2,\tilde{1}}+{\hat{c}^{\dagger}}_{1,\tilde{2}})\hat{Z}^{\dagger}_{\tilde{1}}({\hat{a}^{\dagger}}_{2,\tilde{0}}+{\hat{a}^{\dagger}}_{1,\tilde{1}}+{\hat{a}^{\dagger}}_{0,\tilde{2}})\Big{]}|vac\rangle$ (11)

### v2

@@ Abstract
High-dimensional multipartite entanglement plays a crucial role in quantum information science. However, existing schemes for generating such entanglement become complex and costly as the dimension of quantum units increases. In this work, we overcome the limitation by proposing a significantly enhanced linear optical heralded scheme that generates the $d$-level $N$-partite GHZ state with single-photon sources and linear operations. Our scheme requires $dN$ photons, which is the minimal required photon number, with substantially improved success probability from previous schemes. It employs linear optical logic gates compatible with any qudit encoding system and can generate generalized GHZ states with installments of beamsplitters. With efficient generations of high-dimensional resource states, our work opens avenues for further exploration in high-dimensional quantum information processing.
@@
Numerous advantages arise from entangling multiple quantum units in quantum information processing. The Bell state of qubits facilitates quantum enhancements over classical counterparts [4, 5]. Beyond two-qubit systems, quantum entanglement in multipartite and/or higher-dimensional systems is essential for new quantum technologies such as quantum computing [6, 7, 8, 9, 10], quantum cryptography [11, 12] quantum communication protocols [13, 14, 15, 16].
@@
We contribute to overcoming this limit by providing a significantly enhanced heralded scheme for generating the qudit multipartite entanglement in linear optical networks (LONs). The LON is one of promising systems that can feasibly generate quantum resource states. Non-destructive entanglement generations by heralding are possible in LONs by utilizing ancillary particles to herald the successful operation events without direct detections of the target states. Therefore, photons entangled by heralding become proper resources for quantum computations with multiple gates [20, 21] and also facilitates loophole-free Bell tests [22, 23]. However, the number of additional particles and modes that the heralded schemes require increases as the dimension of entanglement becomes higher, which makes the circuit structures much more intricate. Therefore, it is prerequisite to design optimized circuits that achieve the target state with minimal resources for the efficient high-dimensional quantum information processing.
@@
where $|j\rangle_{k}$ denotes the $k$th qudit being in the $j$th state. The qudit GHZ state reveals intriguing non-classical properties of $N$-partite $d$-level systems such as the GHZ paradox [24, 25, 26, 27] and Bell inequalities [28]. It also becomes a resource state to generate qudit cluster states for the high-dimensional measurement-based quantum computing (MBQC) [29, 30].
Schemes for the qudit GHZ state in various physical setups have been proposed in, e.g., Refs. [31, 32, 33, 29, 34, 35, 36] (see Appendix C for a comparison of these schemes). Particularly, Ref. [29] first suggested a linear optical heralded scheme for $|GHZ_{N}^{d}\rangle$ based on the zero-transmission laws (ZTL) in quantum Fourier transforms (QFTs). However, it suffers from low efficiency and exponentially increasing photon numbers as $N$ and $d$ grows.
In this work, we propose a scheme that generates the GHZ state of any $N$ and $d$ with the most feasible optical elements, i.e., single photons and linear optical operators as in Ref. [29]. Our scheme has several manifest advantages over formerly suggested schemes: First, it is an efficient circuit with $dN$ photons, which is the minimal photon number for generating $|GHZ_{N}^{d}\rangle$ (see Appendix D for the proof) and exponentially higher success probability than that in Ref. [29]. Second, it can easily control the probability amplitudes of each term in the GHZ state by simply placing beamsplitters on the paths of photons, which result in the generation of the generalized qudit GHZ state. Third, it consists of linear optical logic gates that any type of qudit encoding (e.g., optical angular momentum (OAM), multi-rail, time-bin, etc.) can establish.
@@
1. $CN_{d}$: The $d$-level generalized CNOT gate between the spatial mode $j$ ($(\in\{0,1,\cdots,d-1\})$ and the internal state $s$ ($(\in\{0,1,\cdots,d-1\})$, which works on states as
$|s\rangle_{j}={\hat{a}}^{\dagger}_{j,s}|vac\rangle$
$\to$ $~{}CN_{d}|s\rangle_{j}=|s\rangle_{j\oplus_{d}s}={\hat{a}}^{\dagger}_{j\oplus_{d}s,s}|vac\rangle$ (2)
(${\hat{a}^{\dagger}}_{j,s}$ is a photon creation operator, $|vac\rangle$ is the vacuum state, and $\oplus_{d}$ is the addition modular $d$). Its inverse $\overline{CN_{d}}$ works as
$|s\rangle_{j}={\hat{a}}^{\dagger}_{j,s}|vac\rangle$
$\to$ $~{}\overline{CN}_{d}|s\rangle_{j}=|s\rangle_{j\ominus_{d}s}={\hat{a}}^{\dagger}_{j\ominus_{d}s,s}|vac\rangle$ (3)
@@ 2. $F_{d}$: The internal state Fourier-transformation gate,
$F_{d}|s\rangle_{j}=|\tilde{s}\rangle_{j}=\frac{1}{\sqrt{d}}\sum_{r=0}^{d-1}\omega^{sr}|r\rangle_{j}~{}~{}~{}(\omega=e^{i\frac{2\pi}{d}}).$ (4)
@@ 3. $X_{d}$: The $d$-level unitary generalization of $\sigma_{x}$ in qubit,
4. $BS_{t}$: A beam splitter that transmits a photon with transmissivity $t$.
5. $D_{F}$: Number resolving detections in the Fourier-transformed basis, i.e., measurements in $\{|\tilde{s}\rangle\}_{s=0}^{d-1}$ and postselections of the cases when only one particle arrives at the detector. This is achieved by combining $F_{d}$ and $CN_{d}$, i.e.,
@@
By combining the gates listed in Fig. 1, we can design a circuit that generates $|GHZ_{N}^{d}\rangle$ for any $N$ and $d$. Our scheme is given in Fig. 2, which can be considered a linear optical implementation for the sculpting scheme in Ref. [37] to generate the same target state. While the sculpting scheme in Ref. [37] is a conceptual one with the assumption of the existence of high-dimensional deterministic annihilation operators, here we demonstrate that linear optical logic gates can combine to probabilistically generate the qudit GHZ state with heralding. Another crucial advantage of our scheme over Ref. [37] is that the generalized qudit GHZ state can be easily generated by controlling the transmissivity of $BS_{t}$ in Fig. 2. See Appendix B for a more detailed explanation on these points.
In each input mode of the circuit in Fig. 2, we inject $d$ photons with different internal states in the same mode. The operation of each dashed box in the circuit corresponds to the boson subtraction of $(d-1)$ particles, hence $(d-1)N$ subtraction of photons in total. $BS_{t}$s are attached to control the probability amplitude of the target state, hence we can obtain different states in the class of generalized GHZ states by changing the transmissivity of the $BS_{t}$s. It is worth noting that our circuit inherits the cyclic symmetry of the qudit GHZ state.
@@
Fig. 3 enumerates the numerical values of $Log_{10}(P_{suc}(N,d))$ for $d=3,4,5$. Compared to the schemes using the zero-transmission law (ZTL) in Fourier transform interferometers in Ref. [29] (which we call ‘ZTL scheme’ from now on), our scheme is substantially more efficient for $N\geq 3$ both in the photon numbers and success probabilities. Appendix C.2 presents a detailed comparative explanation on the efficiency of two schemes.
Figure 2: The linear optical circuit that generates $|GHZ_{N}^{d}\rangle.$ Beam splitters are attached at the end of the $d$ sub-modes of the first spatial subsystem $1$ to control the probability amplitudes of the terms in the target state, hence we can obtain the generalized qudit GHZ states.
Figure 3: The logarithmic values of $P_{suc}(N,d)$ for $d=3,4,5$ as $N$ increases from 2 to 6. We can see that $P_{suc}(2,3)$ and $P_{suc}(3,3)$ have the order of $10^{-2}$ and $10^{-4}$ respectively. Our scheme is substantially more efficient for $N\geq 3$ than Ref. [29] both in the photon numbers and success probabilities, which is explained in detail in Appendix C.2.
@@
As proof of concept, we provide here a step-by-step analysis on the $N=d=3$ example (Fig. 4). There are three spatial subsystems $\{a,b,c\}$ that has three sub-modes respectively, hence the photon creation operators are denoted as $\{{\hat{a}^{\dagger}}_{j,s},{\hat{b}^{\dagger}}_{k,r},{\hat{c}^{\dagger}}_{l,t}\}$ where $\{j,k,l\}$ denotes the spatial sub-modes and $\{s,r,t\}$ the qudit states ($j,k,l,s,r,t\in\{0,1,2\}$).
@@
Figure 4: $|GHZ_{3}^{3}\rangle$ example of our scheme. We need a BS with transmissivity $\frac{1}{\sqrt{3}}$ on $X$, $Y$ and $Z$ to carve the extra amplitude of $|111\rangle$.
@@
Step 1. Initial state preparation: We inject three photons of orthogonal internal states with each other into the 0th modes of $a$, $b$, and $c$ respectively, i.e.,
${\hat{a}}^{\dagger}_{0,0}{\hat{a}}^{\dagger}_{0,1}{\hat{a}}^{\dagger}_{0,2}{\hat{b}}^{\dagger}_{0,0}{\hat{b}}^{\dagger}_{0,1}{\hat{b}}^{\dagger}_{0,2}\hat{c}^{\dagger}_{0,0}\hat{c}^{\dagger}_{0,1}\hat{c}^{\dagger}_{0,2}|vac\rangle,$ (7)
of which we Fourier-transform the internal state basis:
${\hat{a}}^{\dagger}_{0,\tilde{0}}{\hat{a}}^{\dagger}_{0,\tilde{1}}{\hat{a}}^{\dagger}_{0,\tilde{2}}{\hat{b}}^{\dagger}_{0,\tilde{0}}{\hat{b}}^{\dagger}_{0,\tilde{1}}{\hat{b}}^{\dagger}_{0,\tilde{2}}\hat{c}^{\dagger}_{0,\tilde{0}}\hat{c}^{\dagger}_{0,\tilde{1}}\hat{c}^{\dagger}_{0,\tilde{2}}|vac\rangle.$ (8)
@@
Step 2. $CN_{3}$ operation: This operator works on the basis of $\{|j\rangle\}_{j=0}^{2}$, hence the initial state is transformed as
$\frac{1}{\sqrt{3}^{9}}$ $\big{(}{\hat{a}}^{\dagger 3}_{0,0}+{\hat{a}}^{\dagger 3}_{1,1}+{\hat{a}}^{\dagger 3}_{2,2}-3{\hat{a}}^{\dagger}_{0,0}{\hat{a}}^{\dagger}_{1,1}{\hat{a}}^{\dagger}_{2,2}\big{)}$
×$\times\big{(}{\hat{b}}^{\dagger 3}_{0,0}+{\hat{b}}^{\dagger 3}_{1,1}+{\hat{b}}^{\dagger 3}_{2,2}-3{\hat{b}}^{\dagger}_{0,0}{\hat{b}}^{\dagger}_{1,1}{\hat{b}}^{\dagger}_{2,2}\big{)}$
×$\times\big{(}\hat{c}^{\dagger 3}_{0,0}+\hat{c}^{\dagger 3}_{1,1}+\hat{c}^{\dagger 3}_{2,2}-3\hat{c}^{\dagger}_{0,0}\hat{c}^{\dagger}_{1,1}\hat{c}^{\dagger}_{2,2}\big{)}|vac\rangle$ (9)
@@
Step 3. Rewiring: Now we rewire so that creation operators $\{{\hat{a}^{\dagger}}_{1,1},{\hat{b}^{\dagger}}_{1,1},{\hat{c}^{\dagger}}_{1,1}\}$ are sent to $\{\hat{X}^{\dagger}_{1},\hat{Y}^{\dagger}_{1},\hat{Z}^{\dagger}_{1}\}$ and $\{{\hat{a}^{\dagger}}_{2,2},{\hat{b}^{\dagger}}_{2,2},{\hat{c}^{\dagger}}_{2,2}\}$ to $\{{\hat{b}^{\dagger}}_{2,2},{\hat{c}^{\dagger}}_{2,2},{\hat{a}^{\dagger}}_{2,2}\}$ :
$\frac{1}{\sqrt{3}^{9}}$ $\big{(}{\hat{a}}^{\dagger 3}_{0,0}+\hat{X}^{\dagger 3}_{1}+{\hat{b}}^{\dagger 3}_{2,2}-3{\hat{a}}^{\dagger}_{0,0}\hat{X}^{\dagger}_{1}{\hat{b}}^{\dagger}_{2,2}\big{)}$
×$\times\big{(}{\hat{b}}^{\dagger 3}_{0,0}+\hat{Y}^{\dagger 3}_{1}+\hat{c}^{\dagger 3}_{2,2}-3{\hat{b}}^{\dagger}_{0,0}\hat{Z}^{\dagger}_{1}\hat{c}^{\dagger}_{2,2}\big{)}$
×$\times\big{(}\hat{c}^{\dagger 3}_{0,0}+\hat{Z}^{\dagger 3}_{1}+\hat{a}^{\dagger 3}_{2,2}-3\hat{c}^{\dagger}_{0,0}\hat{X}^{\dagger}_{1}\hat{a}^{\dagger}_{2,2}\big{)}|vac\rangle$ (10)
Step 4. $X_{d}$ and $F_{3}$: $F_{3}$ is applied to both wires in each subsystem, hence the state is transformed to
$\frac{1}{\sqrt{3}^{9}}$ $\big{(}{\hat{a}}^{\dagger 3}_{0,\tilde{0}}+\hat{X}^{\dagger 3}_{1}+{\hat{b}}^{\dagger 3}_{2,\tilde{0}}-3{\hat{a}}^{\dagger}_{0,\tilde{0}}\hat{X}^{\dagger}_{1}{\hat{b}}^{\dagger}_{2,\tilde{0}}\big{)}$
$\times$ $\big{(}{\hat{b}}^{\dagger 3}_{0,\tilde{0}}+\hat{Y}^{\dagger 3}_{1}+{\hat{c}}^{\dagger 3}_{2,\tilde{0}}-3{\hat{b}}^{\dagger}_{0,\tilde{0}}\hat{Y}^{\dagger}_{1}{\hat{c}}^{\dagger}_{2,\tilde{0}}\big{)}$
$\times$ $\big{(}\hat{c}^{\dagger 3}_{0,\tilde{0}}+\hat{Z}^{\dagger 3}_{1}+{\hat{a}}^{\dagger 3}_{2,\tilde{0}}-3\hat{c}^{\dagger}_{0,\tilde{0}}\hat{Z}^{\dagger}_{1}{\hat{a}}^{\dagger}_{2,\tilde{0}}\big{)}|vac\rangle$ (11)
@@
$\frac{1}{3^{9}}\Big{[}({\hat{a}^{\dagger}}_{0,0}+{\hat{a}^{\dagger}}_{2,1}+{\hat{a}^{\dagger}}_{1,2})^{3}+\hat{X}^{\dagger 3}_{1}+({\hat{b}^{\dagger}}_{2,0}+{\hat{b}^{\dagger}}_{1,1}+{\hat{b}^{\dagger}}_{0,2})^{3}$
$~{}~{}~{}-3\sqrt{3}({\hat{a}^{\dagger}}_{0,0}+{\hat{a}^{\dagger}}_{2,1}+{\hat{a}^{\dagger}}_{1,2})\hat{X}^{\dagger}_{1}({\hat{b}^{\dagger}}_{2,0}+{\hat{b}^{\dagger}}_{1,1}+{\hat{b}^{\dagger}}_{0,2})\Big{]}$
×$~{}\times\Big{[}({\hat{b}^{\dagger}}_{0,0}+{\hat{b}^{\dagger}}_{2,1}+{\hat{b}^{\dagger}}_{1,2})^{3}+\hat{Y}^{\dagger 3}_{1}+({\hat{c}^{\dagger}}_{2,0}+{\hat{c}^{\dagger}}_{1,1}+{\hat{c}^{\dagger}}_{0,2})^{3}$
$~{}~{}~{}-3\sqrt{3}({\hat{b}^{\dagger}}_{0,0}+{\hat{b}^{\dagger}}_{2,1}+{\hat{b}^{\dagger}}_{1,2})\hat{Y}^{\dagger}_{1}({\hat{c}^{\dagger}}_{2,0}+{\hat{c}^{\dagger}}_{1,1}+{\hat{c}^{\dagger}}_{0,2})\Big{]}$
×$~{}\times\Big{[}({\hat{c}^{\dagger}}_{0,0}+{\hat{c}^{\dagger}}_{2,1}+{\hat{c}^{\dagger}}_{1,2})^{3}+\hat{Z}^{\dagger 3}_{1}+({\hat{a}^{\dagger}}_{2,0}+{\hat{a}^{\dagger}}_{1,1}+{\hat{a}^{\dagger}}_{0,2})^{3}$
$~{}~{}~{}-3\sqrt{3}({\hat{c}^{\dagger}}_{0,0}+{\hat{c}^{\dagger}}_{2,1}+{\hat{c}^{\dagger}}_{1,2})\hat{Z}^{\dagger}_{1}({\hat{a}^{\dagger}}_{2,0}+{\hat{a}^{\dagger}}_{1,1}+{\hat{a}^{\dagger}}_{0,2})\Big{]}|vac\rangle$ (12)